### v1

# Task-Agnostic Continual Prompt Tuning with Gradient-Based Selection and Decoding

Abstract
Prompt-based continual learning (CL) offers a parameter-efficient way to adapt large language models (LLMs) across task sequences. However, most existing methods assume task-aware inference and maintain a growing list of task-specific prompts, which limits scalability and hides latent forgetting. In this work, we introduce GRID, a unified framework that addresses two key limitations: (1) latent forgetting under task-agnostic inference, and (2) prompt memory explosion as task sequences grow. GRID integrates a task-aware decoding mechanism that improves backward transfer by leveraging representative inputs, automatic task identification, and constrained decoding. Additionally, we propose a gradient-based prompt selection strategy that compresses less informative prompts into a single aggregated representation, enabling scalable and memory-efficient lifelong learning. Extensive experiments across short-sequence, long-sequence, and negative transfer benchmarks show that GRID significantly improves backward transfer, achieves competitive forward transfer, and reduces forgotten tasks by up to 80%, outperforming state-of-the-art methods on T5 and Flan-T5 backbones. Source code is available here.
1 Introduction
Continual learning (CL) [37] enables models to learn from a sequence of tasks without retraining from scratch. CL systems build on prior knowledge while adapting to new tasks, which is crucial in dynamic real-world settings. Recent advancements, especially in NLP [42, 33], have focused on three paradigms: regularization-based methods [20, 17], rehearsal-based methods [28, 36], and architecture-based methods [38, 10]. While rehearsal-based methods are effective, they are impractical in privacy-sensitive scenarios. With the increasing complexity of pretrained models [41], full model finetuning has become infeasible. This has led to the rise of parameter-efficient finetuning (PEFT) techniques [9], such as prompt tuning (PT) [18], which adapts large models by training only a small set of soft prompts, requiring less than 0.01% of the model’s parameters.
Building on this, continual prompt tuning (CPT) extends PT to the CL setting by enabling models to learn task-specific prompts sequentially without modifying the base model [45]. Recent advancements in prompt tuning-based continual learning (CPT) address key challenges like catastrophic forgetting and forward knowledge transfer. [44] introduced a dual prompt framework with a shared and task-specific prompt, though this approach faces limitations in retaining knowledge from past tasks. To mitigate these issues, ProgPrompt [27] maintains a list of prompts for each task, progressively adding new ones while retaining old ones. Coda-prompt [35] extended this by proposing prompt set expansion using weighted summation to ensure efficient knowledge transfer. A substantial body of work has focused on improving CPT’s ability to retain and transfer knowledge, using techniques such as prompt concatenation [27], parameter sharing [45], and weighted prompt summation [15].
Although significant attention has been given, there are still two key challenges in prompt-based CL: (1) latent forgetting under task-agnostic inference and (2) scalability issues from growing prompt queues. First, recent work in prompt-based continuous learning often assumes that assigning a separate prompt to each task eliminates catastrophic forgetting [22], since the model’s core parameters and previously learned prompts remain untouched [18]. Following this assumption, approaches such as Progressive Prompts [27], Q-Tuning [13], and SHLPT [47] report zero Backward Transfer (BWT), suggesting that prior task knowledge is fully preserved. However, these claims are based on task-aware inference, where the identity of the task is known and used to select the corresponding prompt. In real-world scenarios, this assumption often fails: not only is task identity typically unavailable at inference time, but storing a growing number of task-specific prompts becomes increasingly impractical in continual learning settings with an increasing number of tasks. Under such task-agnostic setups, the model must generalize across all previously learned tasks without relying on task-specific prompts. In this setting, our extensive experiments show that the performance on earlier tasks degrades after training on new ones, with the model frequently generating incorrect or ambiguous labels. This issue is particularly problematic for encoder-decoder models like T5, which generate output labels as text. Without clear task guidance, the model may produce label words from unrelated tasks it has seen during pretraining or prior learning. These observations point to a form of latent forgetting that is obscured by task-aware evaluation. In particular, this limitation is also acknowledged in recent works like [13], which find that despite having strong forward transfer, it is not capable of supporting backward knowledge transfer or task-agnostic inference.
The second challenge in prompt-based continual learning lies in addressing the scalability issues in time and memory as the prompt queue grows. Approaches such as Progressive Prompts [27] train and retain a separate soft prompt for each task, resulting in a prompt list that grows linearly with the number of tasks. As task sequences grow, this leads to $\mathcal{O}(N)$ prompt memory and $\mathcal{O}(N^{2})$ inference complexity in transformer-based models. The recent Q-Tuning method [13] addresses this challenge by maintaining a bounded-size prompt queue and discarding less informative prompts using a principal component analysis (PCA)-based eviction strategy. However, it still encounters a trade-off between performance and capacity, and incurs non-negligible computational overhead due to the singular value decomposition (SVD) required for each PCA-based eviction.
In this paper, we aim to address these challenges with the following key contributions:
  • We are the first to reveal that existing prompt-based CL methods reporting $\mathbf{0}$ BWT rely on task-aware inference, which masks latent forgetting in task-agnostic settings. We also identify scalability limitations due to prompt memory growing linearly with task count.
  • We propose GRID (Gradient-based prompt selection with Representative samples selection, task Identification, and Decoding), a unified framework with two core components.
    • A task-aware decoding mechanism that leverages representative inputs and constrained decoding to ensure label consistency and improve backward knowledge retention (BWT) without needing task IDs.
    • A gradient-based prompt selection mechanism that dynamically evaluates the usefulness of prompts and merges less informative ones, reducing storage by 66.7% while maintaining performance.
  • We conduct extensive experiments on encoder-decoder models (T5 and Flan-T5) across short, long, and negative transfer sequences. GRID achieves up to 54% BWT improvement, reduces number of forgotten tasks by 80%, and matches or outperforms state-of-the-art baselines such as ProgPrompt, and SHLPT, under realistic task-agnostic settings.
2 Related Work
Continual Learning. Continual learning (CL) involves learning from a sequence of tasks without full access to previous task data, aiming to preserve prior knowledge and enable positive transfer. A key challenge is catastrophic forgetting [22], where updates to model parameters on new data erode earlier knowledge. CL strategies are categorized into three approaches: memory-based methods store and replay past task data to mitigate forgetting [34, 3]; regularization-based methods penalize deviations from important parameters to retain knowledge without stored data [17, 51]; and architecture-based methods expand the model by adding new components for each task [32, 50]. These methods face scalability issues in large pretrained models, motivating the development of parameter-efficient CL techniques using lightweight components like prompts and adapters [48, 31].
Continual Prompt Tuning. Prompt tuning [18, 19, 12, 43, 14] is a parameter-efficient method for adapting large language models (LLMs) to new tasks by learning a small set of continuous vectors, known as soft prompts, which are prepended to the input tokens. Unlike traditional finetuning, which modifies all model parameters, prompt tuning optimizes only the prompt parameters, leaving the rest of the model unchanged. This approach has been shown to match or surpass the performance of full model finetuning, offering substantial savings in computational cost and memory usage. Building on this, continual prompt tuning [53, 49, 11, 44, 27] extends prompt tuning to the continual learning (CL) setting, where models must adapt to evolving task sequences [18, 12]. However, existing continual prompt tuning methods face challenges, such as the reliance on large memory buffers to prevent forgetting [53, 11], which are impractical in privacy-sensitive or resource-constrained settings. Additionally, task-specific prompts still suffer from latent forgetting in task-agnostic inference scenarios [13], and strategies that incrementally grow prompt representations [27, 44, 35] struggle with scalability. As the number of tasks increases, the size of the prompt grows, leading to efficiency bottlenecks in both training and inference.
Gradient-Based Data and Parameter Selection. Gradient-based strategies have been extensively studied in data selection and coreset construction [2], where the goal is to identify a representative or influential subset of training data [4, 30]. Methods such as Coreset Selection [16, 23] and Gradient Matching [2] leverage the similarity or norm of training gradients to preserve performance while reducing dataset size. Inspired by these ideas, recent works have begun to explore gradient-based criteria for parameter-efficient tuning. In the context of prompt learning, Q-Tuning [13] proposes managing a fixed-size prompt queue but does not use explicit gradient scoring to retain or merge prompts. Our work draws inspiration from the coreset literature and extends it to the prompt space by dynamically evaluating gradient norms of prompt embeddings, allowing us to merge less informative prompts and improve scalability under long task sequences.
3 Background and Motivation
3.1 Prompt-based Continual Learning and Task-Agnostic Inference
Problem Setup. We consider a continual learning (CL) setting in which a model encounters a sequence of $N$ tasks $\mathcal{T}=\{T_{1},T_{2},\dots,T_{N}\}$, where each task $T_{i}$ is associated with a labeled dataset $D_{i}=\{(x_{j},y_{j})\}_{j=1}^{|D_{i}|}$. Here, $x_{j}$ denotes an input instance and $y_{j}\in\mathcal{Y}_{i}$ is the corresponding label. The model is built upon a pretrained encoder-decoder language model $f(\cdot;\theta)$, whose parameters $\theta$ are kept fixed throughout learning. Rather than finetuning $\theta$, we adapt the model to each task $T_{i}$ by learning a soft prompt $\mathbf{p}_{i}\in\mathbb{R}^{l\times d}$, where $l$ denotes the prompt length and $d$ the embedding dimension. After observing tasks $\{T_{1},\dots,T_{t-1}\}$, we maintain a pool of learned prompts $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{t-1}\}$, which serves as a memory of past task adaptations.
When a new task $T_{t}$ arrives, we initialize a new prompt $\mathbf{p}_{t}$ and train it using data from $D_{t}$, concatenated with the existing prompt queue $\mathcal{P}$. The backbone model $f(\cdot;\theta)$ remains frozen during training; only the new prompt $\mathbf{p}_{t}$ is updated. Let $\mathcal{P}^{(t)}=\mathcal{P}\cup\{\mathbf{p}_{t}\}$ denote the prompt configuration used during training for task $T_{t}$. The model prediction is then given by $\hat{y}=f(x;\mathcal{P}^{(t)}).$ The training objective is to minimize: $\mathcal{L}_{t}=\mathbb{E}_{(x,y)\sim D_{t}}[\ell(f(x;\mathcal{P}^{(t)}),y)]$, where $\ell(\cdot)$ is the token-level cross-entropy loss.
Inference Settings. We consider two commonly studied inference modes in continual learning:
1) Task-Aware Inference: The task identity is known at inference time. The model selects the corresponding prompt $\mathbf{p}_{i}$ from the prompt pool $\mathcal{P}$ to perform prediction on task $T_{i}$.
2) Task-Agnostic Inference: The task identity is unknown at inference time. As a result, the model cannot rely on task-specific prompt selection. Instead, it performs prediction using the entire available prompt pool $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{t-1}\}$, which may optionally include an aggregated prompt derived from previous filtering stages. The concatenated prompt sequence is prepended to the input and passed to the model. Formally, the prediction is defined as: $\hat{y}=f(x;\mathcal{P})$.
The task-agnostic evaluation objective then becomes: $\mathcal{L}_{\text{TA}}=\sum_{T_{k}\in\mathcal{T}_{\text{past}}}\mathbb{E}_{(x,y)\sim D_{k}}\left[\ell(f(x;\mathcal{P}),y)\right].$ This formulation evaluates how well the concatenated prompt pool enables generalization to earlier tasks without needing to retrieve or know their corresponding individual prompts. This setting appears in two important continual learning scenarios: 1) Online learning, where a new task arrives and the model must evaluate or detect alignment with previously seen tasks [1]. 2) Retroactive evaluation, where a model trained on a task sequence is later evaluated on older tasks without access to their individual prompts [5, 8].
3.2 Discussion of Challenges in Existing Prompt-based CL
Despite recent progress, prompt-based continual learning methods suffer from critical limitations: Label Drift and Hallucination. In task-agnostic settings, encoder-decoder models generate free-form text outputs without explicit constraints, often resulting in label drift—predictions biased toward labels from unrelated tasks or the pretraining corpus. While the encoder may capture relevant contextual features, the decoder can produce hallucinated outputs that are semantically invalid or inconsistent with the intended task. We provide qualitative examples of such failures in Table 1.
Table 1: Qualitative comparison under task-agnostic inference. Blue indicates label drift, where Progressive Prompts predict semantically related but incorrect labels. Red highlights hallucination, where the model generates invalid outputs not part of the task’s label set.
Task Input Text Label ProgPrompt
BoolQ Did you ever lecture at Harvard? true true</s>
BoolQ The imperialist nation wanted to strangle the economy of its colony. false <pad>Fal
BoolQ The neoclassical canon was rooted in traditional European aesthetics. true <pad>Fal
MNLI Some of the buildings around the city square … colonial period. entailment <pad>True </s>
MNLI The U.S. Army acceded … to keep U.S. forces in place. entailment <pad>4.0
Inconsistent Label Mappings. Prior methods such as Progressive Prompts [27] and SHLPT [47] rely on human-specified mappings (e.g., 0 $\rightarrow$ "negative") to convert numeric labels into text. These mappings are often inconsistent across datasets and introduce semantic misalignment that hinders generalization across tasks. Our Appendix (e.g., Table 21) further highlights the cases where such mappings cause contradictory model behavior.
Lack of Task Awareness. Current methods do not attempt to infer or utilize the underlying task structure. As a result, tasks with similar semantics (e.g., different sentiment classification datasets) are treated independently, missing opportunities for generalization or shared label space alignment. This leads to redundant prompt representations and missed transfer potential.
Unbounded Prompt Growth. Approaches like Progressive Prompts [27] maintain a separate prompt for each task and concatenate them during inference. This causes prompt pool size to grow linearly with the number of tasks, increasing both memory usage and inference time. Moreover, prompt concatenation can unpredictably affect decoding behavior, especially under task-agnostic evaluation.
These theoretical and practical limitations highlight the need for a more structured and scalable framework for prompt-based continual learning under task-agnostic conditions. We provide empirical evidence in Table 1 showing that label hallucination, and semantic drift persists even when the encoder produces reasonable representations. These findings motivate our task-aware input construction and prompt queue management strategies. Furthermore, several open questions surrounding the inference-time behavior of progressive prompts remain unresolved [^1] See discussion at this issue on GitHub., which our work seeks to address.
Figure 1: Overview of the proposed GRID framework. (S1) model receives a stream of tasks with corresponding datasets. (S2) Representative samples are selected via clustering for each task, and task identification is performed to ensure consistent label formats. (S3) Gradient-based prompt selection is applied: prompts from the frozen prompt pool are ranked based on their gradient norms with respect to the current task; (S4) compressed prompt pool is used to train soft prompts for new tasks with the base model frozen. (S5) During inference, constrained decoding ensures predictions are aligned with the identified task semantics.
4 Method
Reformulation. To address the dual challenges of unstable decoding under task-agnostic inference and unbounded prompt pool growth, we propose a unified framework called GRIDGradient-based prompt selection with Representative sample selection, task Identification, and constrained Decoding. GRID integrates two complementary components: 1) A task-aware input pipeline that enhances backward retention and output consistency by selecting representative samples, performing task identification, and applying constrained decoding; 2) A gradient-based prompt scoring mechanism that reduces prompt pool size by identifying and merging less informative prompts while preserving relevant task knowledge. Together, these components enable scalable and reliable continual learning under task-agnostic inference. See Figure 1 for the pipeline and Appendix A for full algorithms.
4.1 Task-Aware Input Construction for Stable Decoding
Encoder-decoder models generate output labels as textual sequences. When task identity is unavailable at inference, unconstrained decoding can lead to label drift, producing labels from unrelated tasks or the pretraining distribution. To mitigate this, we reformulate task-agnostic inference to operate on a refined input space constructed using three components:
1) Representative Input Sampling. To construct a compact yet informative training subset, we select $k=1000$ representative samples per class via clustering, improving upon the random sampling strategy used in prior work [27]. The data set $\mathcal{D}=\{(x_{j},y_{j})\}$ is partitioned by label $y\in\mathcal{Y}$, and the sentences $\mathbf{e}_{j}=f_{\text{embed}}(x_{j})$ are calculated using a pre-training embedding model (all-MiniLM-L6-v2; 29). K-Means clustering is performed within each class to ensure intra-class diversity, and the top $k/C$ samples closest to each cluster center are selected based on cosine similarity: $\text{sim}(\mathbf{e},\mathbf{c})=\frac{\mathbf{e}\cdot\mathbf{c}}{\|\mathbf{e}\|\|\mathbf{c}\|}.$ This yields a balanced subset $\mathcal{D}_{\text{rep}}$ that spans the semantic space of each class. We empirically find that increasing beyond 1k samples per class, the accuracy either saturates or improves negligibly, indicating that larger sample sizes may not result in substantial gains in performance (See Appendix B Table 10), while clustering improves both data efficiency and generalization.
2) Task Identification. In many datasets, task categories are not explicitly provided, and labels often appear in non-descriptive formats (e.g., {0,1} or {choice1, choice2}) that are ambiguous without task context. To resolve this, we implement a hierarchical task identification module that infers the task type $t^{*}$ given a candidate label set $\mathcal{Y}_{i}$ and a sample input $x$. The process begins with rule-based heuristics, which match label tokens and input structure to predefined task templates (e.g., “positive/negative” for sentiment analysis, or premise-hypothesis pairs with “entailment/contradiction” for NLI). If this fails, we fallback to zero-shot task classification using a lightweight generative language model (e.g., Phi-3.5). The model is prompted to infer the task type and remap non-descriptive labels into meaningful textual tokens: $(t^{*},\tilde{\mathcal{Y}}_{i})=\textbf{LLM}(x,\mathcal{Y}_{i}).$ This remapping enables consistent semantic interpretation across tasks, making it possible to apply decoding constraints in a task-agnostic setting.
3) Constrained Decoding. With remapped label set $\mathcal{L}_{i}=\{\ell_{1},\dots,\ell_{K}\}$, we apply constrained decoding at inference. At each decoding step $t$, we restrict the softmax to only allow tokens from $\mathcal{L}_{i}$:
$\tilde{P}(y_{t}\mid y_{<t},x)=\text{softmax}(M\odot\mathbf{z}_{t}),\quad M_{j}=\mathbf{1}[v_{j}\in\mathcal{L}_{i}],$
where $\mathbf{z}_{t}$ are the raw logits and $M$ is a binary mask over the vocabulary $\mathcal{V}$.
4.2 Prompt Pool Compression via Gradient-Based Selection
Prompt-based continual learning methods typically store one prompt per task, leading to prompt pool growth of $\mathcal{O}(N)$. To enable scalable lifelong learning, we dynamically select and merge prompts based on their gradient relevance to the current task. Let $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{t-1}\}$ denote the prompt pool before learning task $T_{t}$. For each prompt $\mathbf{p}_{j}\in\mathcal{P}$, we compute the average gradient norm of the loss $L$ over the new task data $D_{t}$:
$g_{j}=\frac{1}{|D_{t}|}\sum_{(x,y)\in D_{t}}\|\nabla_{\mathbf{p}_{j}}L(f(x;\mathbf{p}_{j}),y)\|_{2}.$ (1)
A high gradient norm indicates that the current task significantly alters the prompt, suggesting it contains distinct knowledge not yet captured—making it valuable for maintaining diversity. In contrast, prompts with low gradient norms align closely with the current task, indicating redundancy; hence, they are compressed to retain the potential for forward transfer while reducing memory footprint. Prompts are partitioned using a threshold:
$\tau=\mu_{g}+\alpha\sigma_{g},$ (2)
where $\mu_{g}$ and $\sigma_{g}$ are the mean and standard deviation of gradient norms, and $\alpha$ is a tunable hyperparameter. We then define the subsets of the prompts as follows:
$\mathcal{P}_{high}=\{\mathbf{p}_{j}\mid g_{j}>\tau\},\quad\mathcal{P}_{low}=\{\mathbf{p}_{j}\mid g_{j}<\tau\}.\quad$ (3)
To prevent the complete loss of information encoded in low-gradient prompts $\mathcal{P}_{low}$, we propose an aggregation mechanism that consolidates them into a single aggregated prompt embedding $\mathbf{p}_{agg}$. We compute this aggregated prompt using gradient-weighted averaging as
$\mathbf{p}_{agg}=\sum_{\mathbf{p}_{j}\in\mathcal{P}_{low}}w_{j}\mathbf{p}_{j},\quad w_{j}=\frac{g_{j}}{\sum_{\mathbf{p}_{k}\in\mathcal{P}_{low}}g_{k}}.$ (4)
This weighting scheme ensures that prompts with relatively higher gradients—even among low-gradient ones—contribute more proportionally to the aggregated prompt. Consequently, the updated prompt pool for task $T_{t}$ is $\mathcal{P}^{\prime}=\mathcal{P}_{high}\cup\{\mathbf{p}_{agg}\}.$ This strategy retains essential knowledge while substantially reducing memory and inference costs. In particular, we observe minimal performance degradation even with a compressed prompt pool, indicating that less relevant prompts offer limited benefit for future tasks.
5 Experiments
5.1 Datasets and Baselines
Datasets. Following Razdaibiedina [27], we evaluate our approach under two few-shot continual-learning settings: the “short-sequence” experiments and the “long-sequence” experiments. In addition, we introduce a Negative Transfer benchmark composed of sequences specifically chosen to induce transfer degradation, to assess robustness when task similarity is low. We also construct three long-sequence benchmark (order A-C) to analyze performance across different task difficulty progressions.
Long-Sequence CL Setting. We propose three novel task orders: (1) a Order $A$: easy-to-hard order, where tasks are categorically grouped and simpler tasks such as binary sentiment classification are introduced first, followed by more complex reasoning tasks like MultiRC and COPA; (2) a Order $B$: hard-to-easy order that reverses this progression; and (3) a Order $C$: mixed order that randomly interleaves tasks from different categories and difficulties. These variations allow us to assess how the order of task presentation and semantic similarity affect forgetting, transfer, and model generalization in long-horizon continual learning scenarios.
Negative Transfer Benchmark. To evaluate the robustness of our approach against negative transfer [24], we introduce a Negative Transfer Benchmark (NT1, NT2, NT3) inspired by the SHLPT framework [47]. Specifically, we identify task pairs that exhibit negative transfer based on the analysis provided in the SHLPT. Using these insights, we create three task sequences that are likely to cause negative transfer, meaning earlier tasks make it harder to learn later ones. This helps us test how well our method handles such challenges.
Baseline Methods for Comparison. In our experiments, we evaluate the proposed method against a total of 11 baseline methods222we could not compare with Q-Tuning[13] as they adopt a different setup and their codebase is not provided.: (1) Finetune [46], (2) Continual Finetune [46], (3) Prompt Tuning [18, 25], (4) Data Replay [8], (5) EWC [17], (6) A-GEM [6], (7) LFPT5 [25], (8) Per-task Prompt [18, 25], (9) ProgPrompt [27], (10) Coda-prompt [35], and (11) SHLPT [47].
We evaluate our method using multiple continual learning benchmarks. Specifically, we present the average test accuracy, backward transfer (BWT), and forward transfer (FWT) scores across different benchmark settings: short-sequence, long-sequence, easy-hard mixed, and negative transfer benchmark. All results are averaged over three runs for robustness. Further experimental metrics and details are provided in Appendix B.
(A) Baseline
(B) Ours
(C) B - A
Figure 2: Heatmaps of backward transfer scores on previous tasks for Order 4. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
5.2 Backward Transfer Analysis
Table 3 presents the Backward Transfer (BWT) scores for long sequence, measuring how much the model forgets previous tasks while learning new ones. Our approach consistently reduces forgetting across all model variant compared to the baselines. Interestingly, although larger models generally achieve higher average accuracy (as seen in Table 12 in the Appendix), they tend to exhibit greater forgetting. In contrast, smaller models achieve better (i.e., less negative) BWT scores. One plausible explanation is that smaller models, with fewer parameters, may learn less aggressively on the current task, thereby generalizing better and retaining prior knowledge more effectively. This highlights a nuanced trade-off between model capacity and forgetting behavior in continual learning.
We observe consistent trends under the negative transfer benchmarks (NT1–NT3). As shown in Table 14 (Appendix), our method significantly reduces forgetting across all model sizes. Notably, Smaller models again show superior BWT, highlighting their robustness to task dissimilarity and the overall effectiveness of our approach in mitigating forgetting.
To visualize BWT dynamics, Figures 2 and 3 offer detailed breakdowns. Figure 2 displays heatmaps of BWT scores for Order 4. Compared to the baseline (A), our method (B) exhibits brighter regions in the lower triangle, indicating stronger retention of earlier tasks. The difference heatmap (C) highlights widespread positive gains (blue), especially for early tasks, underscoring the effectiveness of compressing redundant prompts while preserving diverse, task-relevant knowledge. Figure 3 shows per-task BWT differences between our method and the baseline. Positive values indicate improved retention, while negative values signify degradation. Notably, our method yields significant BWT improvements on challenging tasks like boolq, sst2, and dbpedia_14, while maintaining parity or improvement on most others. This further confirms the effectiveness of our gradient-based compression in preserving essential task knowledge over long sequences.
One critical observation from our results is that Order B consistently outperforms Order A in both BWT scores and the number of forgotten tasks. Specifically, for the same set of tasks, Order B shows less negative BWT scores, indicating reduced forgetting. Additionally, the number of forgotten tasks is lower in Order B, further suggesting that harder tasks at the beginning of training help the model retain earlier knowledge more effectively. This observation highlights the benefit of starting with more difficult tasks, as it leads to better long-term retention and improved overall performance in continual learning. One also notes that Order C performs better than Order A but slightly less effectively than Order B, indicating that a mixed task sequence provides a balanced approach to task progression, offering a compromise between the benefits of easy-to-hard and hard-to-easy learning.
Figure 3: Per-task BWT comparison between our method (blue) and the baseline (red) for Order 4. Positive bars indicate improved retention of prior tasks. Our method shows significant BWT gains on several tasks (e.g., copa, wic, yahoo), demonstrating its effectiveness in mitigating forgetting across diverse task types.
Forgotten Task Count. We further analyze the number of forgotten tasks across various settings in Table 2. Our method significantly reduces the number of forgotten tasks compared to the Progressive Prompts baseline. For instance, in the case of T5-small, the number of forgotten tasks drops from as high as 93 in some orders to as low as 7 under our method. Similar improvements are observed for T5-base, where the count decreases from 90 to just 6 in the best case, and for T5-large, where it reduces from 87 to only 5. These improvements are consistent across all task orderings, including structured sequences such as Order A (easy-to-hard), Order B (hard-to-easy), and Order C (mixed). The results demonstrate that our gradient-based prompt compression effectively mitigates catastrophic forgetting, particularly in long-horizon continual learning setups, while preserving task-relevant knowledge in a compact form.
Table 2: Comparison of forgotten task counts between Progressive Prompts (Baseline) and our method for different models across various task orders (Order 4–6, A–C), now including mean forgotten tasks and relative improvement.
Model Setting Order 4 Order 5 Order 6 Order A Order B Order C Average Improvement (%)
T5-small ProgPrompt 85 85 88 90 84 93 87.5 76.8%
Ours 14 7 28 36 16 21 20.3
FT5-small ProgPrompt 93 95 95 94 97 93 94.5 76.7%
Ours 17 14 28 26 20 27 22.0
T5-base ProgPrompt 88 82 76 90 83 88 84.5 80.9%
Ours 15 6 19 29 13 15 16.2
FT5-base ProgPrompt 94 93 93 93 93 93 93.2 83.0%
Ours 10 10 15 25 14 21 15.8
T5-large ProgPrompt 77 72 80 87 71 85 78.7 82.4%
Ours 11 5 18 26 10 13 13.8
FT5-large ProgPrompt 92 93 93 93 91 93 92.5 79.8%
Ours 15 10 24 24 14 25 18.7
Table 3: Backward Transfer (BWT) comparison between our method and ProgPrompt across multiple task orderings (Order 4–6, A–C) for T5 and Flan-T5 models. Less negative scores indicate reduced forgetting. Our method consistently improves BWT across all model variants.
Model Setting Order 4 Order 5 Order 6 Order A Order B Order C Avg Imp. (%)
T5-small ProgPrompt -0.6081 -0.5414 -0.5126 -0.5243 -0.5377 -0.5006 -0.5374 54.4 $\uparrow$
Ours -0.1402 -0.1894 -0.2504 -0.3141 -0.2734 -0.3006 -0.2447
FT5-small ProgPrompt -0.5711 -0.5574 -0.4916 -0.4914 -0.5075 -0.4987 -0.5196 68.4 $\uparrow$
Ours -0.0518 -0.1449 -0.2106 -0.2129 -0.1517 -0.2396 -0.1686
T5-base ProgPrompt -0.6918 -0.6520 -0.5948 -0.5956 -0.6471 -0.5881 -0.6282 52.7 $\uparrow$
Ours -0.3060 -0.2404 -0.3549 -0.2893 -0.3439 -0.3259 -0.3101
FT5-base ProgPrompt -0.6092 -0.7222 -0.6488 -0.6273 -0.6631 -0.6225 -0.6489 58.7 $\uparrow$
Ours -0.2616 -0.3243 -0.3319 -0.3245 -0.3240 -0.2743 -0.3067
T5-large ProgPrompt -0.7275 -0.7625 -0.6137 -0.6257 -0.6351 -0.6416 -0.6677 51.7 $\uparrow$
Ours -0.3243 -0.3336 -0.3979 -0.3912 -0.2956 -0.3512 -0.3490
FT5-large ProgPrompt -0.7679 -0.7444 -0.6195 -0.6616 -0.7107 -0.6631 -0.6945 50.4 $\uparrow$
Ours -0.3115 -0.3656 -0.3544 -0.3514 -0.3282 -0.3540 -0.3442
5.3 Forward Transfer Analysis
We also report FWT scores in Appendix B, showing that our improvements in BWT do not come at the cost of FWT. In fact, our method achieves comparable or even superior FWT in several cases, indicating that the prompt selection strategy not only mitigates forgetting but also avoids incorporating suboptimal prompts that could harm current task performance. These results highlight that enhancing BWT can be achieved without compromising forward transfer when prompts are selected judiciously.
5.4 Average Accuracy
Table 4: Average test-set accuracy on long-sequence experiments with the T5-large model.
Method T5-large
Order 4 Order 5 Order 6 Avg
Finetune 7.4 7.4 7.5 7.4
Prompt Tuning 8.2 8.7 7.9 10.9
Data Replay 55.2 54.8 53.3 54.4
LFPT5 70.4 68.2 69.1 69.2
Per-task Prompt 77.0 77.0 77.0 77.0
CODA-Prompt 76.2 77.0 76.4 76.5
ProgPrompt 75.7 78.6 74.3 76.2
SHLPT 77.6 76.9 77.8 77.4
Ours* 79.1 80.7 81.1 80.3
On the short-sequence setup (Table 20 in the Appendix), it attains 74.60% accuracy with T5-large, slightly outperforming ProgPrompt (74.5%) and surpassing baselines like LFPT5 and Continual Finetune. Since the short-sequence setup involves only four tasks, we retain all prompt candidates without compression. However, CODA-Prompt and SHLPT outperform our method slightly. These marginal differences are less significant given the simpler nature of short-sequence tasks. In contrast, our method excels in the long-sequence experiments (Table 4), achieving an average accuracy of 80.3%, surpassing all other baselines, including CODA-Prompt and SHLPT, demonstrating its superior performance in handling more complex, long-term task sequences. These results validate the effectiveness of our gradient-driven pruning approach in reducing redundancy while preserving essential task knowledge over long sequences. Detailed results across T5 and Flan-T5 model variants (Table 12 in the Appendix) show consistent gains, particularly for larger models (e.g., +3.2% for T5-large, +5.2% for FT5-base), highlighting the scalability and robustness of our method. Under the Negative Transfer benchmark (Orders NT1–NT3; Table 14 in the Appendix), our approach again outperforms all baselines, with improvements such as +3.8% on FT5-base and +3.6% on FT5-large, demonstrating strong generalization even in low-task-similarity scenarios.
5.5 Ablation Study
Table 5: Ablation study on the GRID framework over Orders 4–6, evaluating the contribution of gradient-based prompt selection (G) and task-aware decoding (D).
Model Variants 4 5 6 Avg
T5-large (0) G.R.I.D. -0.3243 -0.3336 -0.3979 –0.3519
(1) (w/o G) -0.3254 -0.3321 -0.3895 –0.3490
(1)+(2) w/o (D) -0.7032 -0.7589 -0.5967 –0.6863
(3) w/o G.R.I.D -0.7275 -0.7625 -0.6137 –0.7012
FT5-large (0) G.R.I.D. -0.3115 -0.3656 -0.3705 –0.3492
(1) (w/o G) -0.3145 -0.3510 -0.3723 –0.3459
(1)+(2) w/o (D) -0.7321 -0.7428 -0.6056 –0.6935
(3) w/o G.R.I.D -0.7679 -0.7444 -0.6195 –0.7106
To assess the contribution of each component in the GRID framework, we perform an ablation study on task Orders 4, 5 and 6. We evaluate the BWT scores under four configurations: (0) the complete G.R.I.D. model, (1) removing the gradient-based prompt selection (G), (2) removing task-aware decoding (D), and (3) removing both components (i.e., no GRID). As shown in Table 5, the full GRID model (row 0) achieves the best average BWT scores for both T5-large (-0.3519) and FlanT5-large (-0.3492), indicating strong resistance to forgetting. When only the gradient-based selection is removed (row 1), the performance remains nearly unchanged. The comparable scores confirm that prompt compression via gradient selection does not degrade retention. This is expected, as the goal of the gradient-based strategy is not to improve BWT directly, but to reduce prompt memory usage while maintaining similar performance.
Table 6: Prompt memory usage.
Method KB Reduction
ProgPrompt 600
Ours 200 66.7%
This memory advantage is quantitatively shown in Table 6, where our method requires only 200 KB of storage compared to 600 KB for ProgPrompt, achieving a 66.7% reduction. In contrast, removing the task-aware decoding component (row 2) leads to a noticeable drop in BWT, and removing both components (row 3) results in the worst performance for both models (-0.7012 for T5-large, -0.7106 for FlanT5-large). This highlights the critical role of decoding in backward knowledge retention, and the synergy between decoding and prompt selection in the full GRID setup. Overall, this ablation confirms that while the decoding component is central to improving BWT, the gradient-based prompt selection plays a complementary role by enabling scalability without compromising performance.
Impact of Prompt Selection Strategies. Table 7 compares prompt selection strategies (Random, FIFO, and Gradient-based) on T5-large and FlanT5-large for Orders 4–6. Our gradient-based method consistently achieves the lowest forgotten task counts and competitive or best BWT scores across all settings. While FIFO occasionally matches our BWT, it retains more redundant prompts, leading to higher forgetting. Random selection performs worst overall, with unstable BWT and the highest forgetting. These results highlight the effectiveness of our approach in reducing memory while preserving prior knowledge.
Table 7: Comparison of prompt selection strategies (Random, FIFO, Gradient-based) for T5-large and FlanT5-large across Orders 4–6. Metrics: average accuracy (Acc), BWT, and number of forgotten tasks (FT).
Model Strategy Order 4 Order 5 Order 6
Acc BWT FT Acc BWT FT Acc BWT FT
T5-large Random 76.81 -0.3482 15 82.07 -0.3549 7 82.06 -0.4065 21
FIFO 76.08 -0.3389 14 57.39 -0.3780 11 81.34 -0.3962 20
Ours 79.12 -0.3243 11 80.69 -0.3336 5 81.09 -0.3979 18
FlanT5-large Random 78.75 -0.3767 18 76.15 -0.3835 12 76.28 -0.3705 25
FIFO 76.62 -0.3481 16 77.05 -0.3981 16 76.60 -0.3633 24
Ours 79.77 -0.3115 15 78.25 -0.3656 10 76.09 -0.3544 24
6 Conclusion
We proposed GRID, a unified framework for scalable and forgetting-resilient prompt-based continual learning. GRID addresses two critical challenges in continual prompt tuning: latent forgetting under task-agnostic inference and unbounded prompt growth as task sequences increase. By combining task-aware decoding with a novel gradient-based prompt selection and compression strategy, GRID enables efficient prompt reuse and effective task retention. Our method significantly improves backward knowledge retention, reducing the number of forgotten tasks by up to 91.8% in certain settings, while maintaining competitive forward transfer. Extensive experiments across short-sequence, long-sequence, and negative transfer benchmarks, using both T5 and Flan-T5 backbones, demonstrate that GRID achieves state-of-the-art performance in both accuracy and forgetting metrics, making it a practical solution for realistic, long-horizon continual learning.
Appendix A Algorithmic Details
Algorithm 1 Representative Sample Selection
1:Dataset $\mathcal{D}=\{(x_{j},y_{j})\}$, samples per class $k$, embedding model $f_{\text{embed}}$, number of clusters $C$
2:Subset $\mathcal{D}_{\text{rep}}$
3:Initialize $\mathcal{D}_{\text{rep}}\leftarrow\emptyset$
4:for each label $y$ do
5:    $\mathcal{D}_{y}\leftarrow\{x_{j}\mid y_{j}=y\}$
6:    Embed each $x_{j}$: $\mathbf{e}_{j}\leftarrow f_{\text{embed}}(x_{j})$
7:    Run K-Means on $\{\mathbf{e}_{j}\}$ into $C$ clusters
8:    for each cluster $c$ do
9:        Select top $k/C$ samples by similarity to center
10:        Add to $\mathcal{D}_{\text{rep}}$
11:    end for
12:    if selected < $k$ then
13:        Add random samples to reach $k$ total
14:    end if
15:end for
16:Shuffle and return $\mathcal{D}_{\text{rep}}$
Algorithm 2 Prompt Selection and Aggregation
1:Prompt pool $\mathcal{P}=\{\mathbf{p}_{j}\}$, new task data $D_{t}$, thresholds $\alpha,\beta$
2:Updated prompt pool $\mathcal{P}^{\prime}$
3:for each $\mathbf{p}_{j}\in\mathcal{P}$ do
4:    Compute gradient norm $g_{j}$ via Eq. 1
5:end for
6:Compute $\mu_{g}$, $\sigma_{g}$; threshold $\tau$ via Eq. 2
7:Classify prompts into $\mathcal{P}_{high}$ and $\mathcal{P}_{low}$ via Eq. 3
8:if $\mathcal{P}_{low}\neq\emptyset$ then
9:    Compute aggregation $\mathbf{p}_{\text{agg}}$ via Eq. 4
10:    $\mathcal{P}^{\prime}\leftarrow\mathcal{P}_{high}\cup\{\mathbf{p}_{\text{agg}}\}$
11:else
12:    $\mathcal{P}^{\prime}\leftarrow\mathcal{P}_{high}$
13:end if
14:Append new prompt $\mathbf{p}_{t}$ to $\mathcal{P}^{\prime}$
15:return $\mathcal{P}^{\prime}$
Appendix B Further Implementation Details and Additonal Results
B.1 Datasets
Table 8: Overview of the 15 datasets used in our CL experiments, including their evaluation metrics. Datasets from CL benchmark [52], GLUE [39], and SuperGLUE [40] were utilized, along with the IMDB movie reviews dataset. For tasks with two evaluation metrics, we report the average as the final performance measure.
Dataset name Category Task Domain Metric
1. Yelp CL benchmark sentiment analysis Yelp reviews accuracy
2. Amazon CL benchmark sentiment analysis Amazon reviews accuracy
3. DBpedia CL benchmark topic classification Wikipedia accuracy
4. Yahoo CL benchmark QA Yahoo Q&A accuracy
5. AG News CL benchmark topic classification news accuracy
6. MNLI GLUE NLI various accuracy
7. QQP GLUE paraphrase detection Quora accuracy & F1
8. RTE GLUE NLI news, Wikipedia accuracy
9. SST2 GLUE sentiment analysis movie reviews accuracy
10. WiC SuperGLUE word sense disambiguation lexical databases accuracy
11. CB SuperGLUE NLI various accuracy
12. COPA SuperGLUE QA blogs, encyclopedia accuracy
13. BoolQ SuperGLUE boolean QA Wikipedia accuracy
14. MultiRC SuperGLUE QA various F1 & EM
15. IMDB Other sentiment analysis movie reviews accuracy
Short-Sequence CL Setting:
We evaluate our method on four widely-used text-classification datasets introduced by [52]: Amazon Reviews, DBpedia, Yahoo Answers, and AG News. We follow the few-shot continual learning setting outlined by [25] and [27], evaluating across three distinct task sequences (Orders 1–3) [27].
Long-Sequence CL Setting:
Following the approach of [27], we consider a total of 15 distinct tasks. These tasks include the four previously described in the short-sequence CL benchmark, four tasks from the GLUE benchmark (MNLI, QQP, RTE, SST2) [39], five tasks from the SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQ) [40], as well as YELP reviews [52] and the IMDB movie reviews dataset [21]. We evaluate our methods across six different task sequence orders. Three of these sequences follow standard long-sequence continual learning orders that have been used in prior works (e.g., Orders 4–6)[27, 13]. We additionally propose three novel task orders: (1) a Order $A$: easy-to-hard order, where tasks are categorically grouped and simpler tasks such as binary sentiment classification are introduced first, followed by more complex reasoning tasks like MultiRC and COPA; (2) a Order $B$: hard-to-easy order that reverses this progression; and (3) a Order $C$: mixed order that randomly interleaves tasks from different categories and difficulties. These variations allow us to assess how the order of task presentation and semantic similarity affect forgetting, transfer, and model generalization in long-horizon continual learning scenarios.
B.2 Limitations and Future Work
While GRID demonstrates strong improvements in backward transfer and scalability, it has two primary limitations. First, our evaluation is limited to sequences of 15 tasks. To fully assess the scalability of our method, future work should test it on significantly longer task streams. Second, the current GRID framework—specifically the task-aware decoding component—is designed for encoder-decoder architectures. Although our gradient-based prompt selection strategy is model-agnostic and can be applied broadly, the full framework has not yet been extended to encoder-only or decoder-only models. Exploring how to adapt GRID’s decoding mechanism to these architectures is a promising direction for future research.
Table 9: Task sequence orders used in our CL experiments. Orders 1-3 represent standard short-sequence CL benchmarks, while Orders 4-6 correspond to long-sequence CL benchmarks. Orders A-C are our proposed custom sequences, with A representing an easy-to-hard order, B representing a hard-to-easy order, and C representing a mixed order. NT1, NT2, and NT3 represent different task evaluation orders designed to examine negative transfer across various task sequences.
Order Task Sequence
1 db $\rightarrow$ amazon $\rightarrow$ yahoo $\rightarrow$ ag
2 db $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ yahoo
3 yahoo $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ db
4 mnli $\rightarrow$ cb $\rightarrow$ wic $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ boolq $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ sst2 $\rightarrow$ dpedia $\rightarrow$ ag $\rightarrow$ multirc $\rightarrow$ yahoo
5 multirc $\rightarrow$ boolq $\rightarrow$ wic $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ sst2 $\rightarrow$ dpedia $\rightarrow$ ag $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ yahoo
6 yelp $\rightarrow$ amazon $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ sst2 $\rightarrow$ dpedia $\rightarrow$ ag $\rightarrow$ yahoo $\rightarrow$ multirc $\rightarrow$ boolq $\rightarrow$ wic
A sst2 $\rightarrow$ imdb $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ yahoo $\rightarrow$ dbpedia $\rightarrow$ mnli $\rightarrow$ rte $\rightarrow$ cb $\rightarrow$ qqp $\rightarrow$ copa $\rightarrow$ boolq $\rightarrow$ wic $\rightarrow$ multirc
B multirc $\rightarrow$ wic $\rightarrow$ boolq $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ cb $\rightarrow$ rte $\rightarrow$ mnli $\rightarrow$ dbpedia $\rightarrow$ yahoo $\rightarrow$ ag $\rightarrow$ amazon $\rightarrow$ yelp $\rightarrow$ imdb $\rightarrow$ sst2
C sst2 $\rightarrow$ copa $\rightarrow$ ag $\rightarrow$ imdb $\rightarrow$ mnli $\rightarrow$ yahoo $\rightarrow$ rte $\rightarrow$ yelp $\rightarrow$ qqp $\rightarrow$ cb $\rightarrow$ amazon $\rightarrow$ dbpedia $\rightarrow$ boolq $\rightarrow$ wic $\rightarrow$ multirc
NT1 multirc $\rightarrow$ wic $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ rte $\rightarrow$ qqp $\rightarrow$ yahoo $\rightarrow$ yelp $\rightarrow$ amazon
NT2 amazon $\rightarrow$ yelp $\rightarrow$ yahoo $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ cb $\rightarrow$ mnli $\rightarrow$ wic $\rightarrow$ multirc
NT3 yahoo $\rightarrow$ mnli $\rightarrow$ amazon $\rightarrow$ cb $\rightarrow$ yelp $\rightarrow$ rte $\rightarrow$ qqp $\rightarrow$ multirc $\rightarrow$ wic
Table 10: Accuracy for different datasets with varying sample sizes.
Dataset No. of samples/k Accuracy
dbpedia 20 0.5203
200 0.9674
1k 0.9880
1.5k 0.9892
amazon 20 0.0000
200 0.0000
1k 0.5136
1.5k 0.5387
ag 20 0.0000
200 0.8180
1k 0.8900
1.5k 0.8933
Table 11: Impact of prompt length on T5-large performance across Orders 4–6. We observe a trade-off: longer prompts improve BWT (less forgetting) but increase the number of forgotten tasks and can degrade accuracy. Optimal performance is generally observed at a moderate prompt length (e.g., 10).
Prompt Length Order 4 Order 5 Order 6
Acc BWT FT Acc BWT FT Acc BWT FT
5 78.23 -0.3167 10 81.73 -0.3565 5 84.98 -0.4286 15
10 79.12 -0.3243 11 80.69 -0.3336 5 81.09 -0.3979 18
20 80.58 -0.3396 12 79.37 -0.3159 5 75.55 -0.3206 24
Table 12: Average test accuracy across long-sequence task orderings (Order 4–6, A–C) for T5 and Flan-T5 models. Our method consistently outperforms Progressive Prompts in nearly all configurations.
Model Setting Order 4 Order 5 Order 6 Order A Order B Order C Avg
T5-small ProgPrompt 63.18 57.18 61.12 62.73 59.11 60.33 60.61
Ours 59.75 64.97 59.90 64.64 64.51 62.39 62.69
FT5-small ProgPrompt 60.32 59.25 59.43 59.18 59.35 59.96 59.58
Ours 58.35 57.37 64.73 61.74 58.30 57.07 59.59
T5-base ProgPrompt 71.34 67.60 69.57 69.71 72.43 69.93 70.10
Ours 72.74 71.24 72.09 67.16 75.72 65.23 70.70
FT5-base ProgPrompt 63.73 74.83 75.23 72.14 74.25 72.48 72.11
Ours 77.24 78.08 77.16 78.32 73.53 65.92 75.04
T5-large ProgPrompt 75.68 78.56 74.29 75.05 77.10 75.46 76.02
Ours 79.12 80.69 81.09 79.05 79.79 75.54 79.21
FT5-large ProgPrompt 79.81 77.51 74.30 79.13 79.09 79.56 78.23
Ours 79.77 78.25 76.09 76.91 76.94 76.74 77.45
Table 13: Backward Transfer (BWT) scores on negative transfer benchmarks (NT1–NT3) comparing our method with Progressive Prompts across T5 and Flan-T5 models. Lower-magnitude (less negative) values indicate reduced forgetting. Our method consistently achieves better BWT, demonstrating robustness under task dissimilarity.
Model Setting NT1 NT2 NT3 Avg
T5-small Baseline -0.5783 -0.4882 -0.5133 -0.5266
Ours -0.1957 -0.1945 -0.2043 -0.1982
FT5-small Baseline -0.4771 -0.3924 -0.4097 -0.4264
Ours -0.0807 -0.1601 -0.1171 -0.1193
T5-base Baseline -0.5707 -0.5834 -0.5890 -0.5810
Ours -0.1817 -0.2103 -0.2314 -0.2078
FT5-base Baseline -0.6719 -0.5335 -0.6116 -0.6057
Ours -0.2911 -0.2074 -0.3162 -0.2716
T5-large Baseline -0.7088 -0.6598 -0.6536 -0.6741
Ours -0.3154 -0.3658 -0.3077 -0.3296
FT5-large Baseline -0.7156 -0.6628 -0.6886 -0.6890
Ours -0.2858 -0.2888 -0.3636 -0.3127
Table 14: Average test accuracy on negative transfer task sequences (NT1–NT3) comparing our method with Progressive Prompts. Our method consistently outperforms the baseline across all models, demonstrating better generalization under reduced task similarity.
Model Setting NT1 NT2 NT3 Avg
T5-small Baseline 57.22 55.51 57.79 56.84
Ours 58.21 55.08 56.44 56.58
FT5-small Baseline 51.57 46.02 48.36 48.65
Ours 50.40 50.54 49.77 50.24
T5-base Baseline 61.18 65.75 66.42 64.45
Ours 63.34 63.46 63.30 63.37
FT5-base Baseline 70.81 60.83 67.25 66.30
Ours 70.76 61.54 70.05 67.45
T5-large Baseline 73.57 73.63 74.26 73.82
Ours 72.29 74.84 73.07 73.40
FT5-large Baseline 75.98 73.98 74.84 74.93
Ours 76.28 71.56 74.72 74.19
Table 15: Backward Transfer (BWT) scores on general task orderings (Order 1–3) comparing our method with Progressive Prompts across multiple models. Our method consistently achieves less negative BWT, indicating improved retention of prior tasks across varying task sequences.
Model Setting Order 1 Order 2 Order 3 Avg
T5-small Baseline -0.5284 -0.5214 -0.5899 -0.5466
Ours -0.3102 -0.3010 -0.2471 -0.2861
FT5-small Baseline -0.6116 -0.5871 -0.6462 -0.6150
Ours -0.3590 -0.3254 -0.3265 -0.3370
T5-base Baseline -0.6499 -0.6080 -0.6406 -0.6328
Ours -0.4039 -0.3494 -0.2755 -0.3430
FT5-base Baseline -0.6594 -0.6457 -0.6991 -0.6681
Ours -0.4187 -0.3722 -0.2679 -0.3529
T5-large Baseline -0.6407 -0.6311 -0.7081 -0.6600
Ours -0.4221 -0.3705 -0.3193 -0.3706
FT5-large Baseline -0.6808 -0.6797 -0.7510 -0.7038
Ours -0.4863 -0.4099 -0.3345 -0.4103
Table 16: Average test accuracy on short-sequence continual learning experiments (Order 1–3). Our method performs competitively across all model variants, matching or outperforming the baseline, especially for larger models like FT5-base, while not applying prompt compression due to the limited number of tasks.
Model Setting Order 1 Order 2 Order 3 Avg
T5-small Baseline 65.29 66.11 67.55 66.32
Ours 64.43 64.94 66.05 65.14
FT5-small Baseline 72.15 70.04 72.75 71.65
Ours 72.79 70.90 72.81 72.17
T5-base Baseline 75.70 73.30 72.55 73.85
Ours 75.25 75.65 74.56 75.15
FT5-base Baseline 76.87 76.92 77.12 76.97
Ours 77.59 77.56 76.09 77.08
T5-large Baseline 75.13 77.32 77.73 76.73
Ours 74.84 74.41 74.55 74.60
FT5-large Baseline 80.40 81.23 81.20 80.94
Ours 80.29 80.95 80.61 80.62
Table 17: Forward Transfer (FWT) scores across various task orderings (Order 4–6, A–C) for T5 and Flan-T5 models. Higher scores indicate better transfer to future tasks. Our method achieves comparable or improved FWT in most cases, demonstrating that reducing forgetting does not come at the cost of forward transfer.
Model Setting Order 4 Order 5 Order 6 Order A Order B Order C Avg
T5-small Baseline -5.07 -11.28 -1.48 -6.09 -9.58 -8.63 -7.02
Ours -5.82 -0.11 -5.55 -0.96 -1.11 -3.33 -2.81
FT5-small Baseline -3.57 -5.26 -4.98 -5.33 -4.92 -4.64 -4.78
Ours -2.53 -4.09 3.33 0.25 -3.02 -4.64 -1.78
T5-base Baseline -3.02 -7.02 -4.80 -4.55 -1.58 -4.23 -4.20
Ours -4.93 -6.68 -5.78 -11.32 -2.21 -13.28 -7.37
FT5-base Baseline -13.91 -2.08 -1.48 -4.90 -2.60 -4.53 -4.92
Ours 4.27 5.10 4.04 5.32 0.18 -7.90 1.84
T5-large Baseline -4.14 -1.19 -5.58 -2.66 -2.67 -3.81 -3.34
Ours -2.82 -1.20 -0.90 -3.11 -2.20 -6.28 -2.75
FT5-large Baseline -0.73 -3.17 -6.65 -1.43 -1.53 -0.99 -2.42
Ours 2.35 0.75 -1.48 -0.64 -0.60 -0.90 -0.09
Table 18: Comparison of FWT scores between Progressive Prompts (Baseline) and Ours across different models. The table reports average performance on different tasks order , including Order NT1, NT2, NT3, and the overall average (Avg).
Model Setting NT1 NT2 NT3 Avg
T5-small Baseline -5.86 -7.62 -5.62 -6.37
Ours -4.67 -7.65 -6.87 -6.40
FT5-small Baseline -3.93 -9.99 -7.88 -7.27
Ours -4.96 -4.58 -6.47 -5.34
T5-base Baseline -7.44 -1.35 -1.44 -3.41
Ours -6.52 -5.56 -6.64 -6.24
FT5-base Baseline -1.29 -12.68 -5.35 -6.44
Ours -1.43 -11.72 -2.05 -5.07
T5-large Baseline -2.55 -2.34 -1.62 -2.17
Ours -4.70 -1.85 -3.76 -3.44
FT5-large Baseline 0.09 -2.19 -0.76 -0.95
Ours 1.06 -4.39 -0.77 -1.37
Table 19: Forward Transfer (FWT) scores on short-sequence tasks (Orders 1–3) comparing our method with Progressive Prompts for T5-small and FT5-small. While our method prioritizes backward retention, it maintains competitive FWT performance, confirming that prompt compression does not significantly hinder future task generalization.
Model Setting Order 1 Order 2 Order 3 Avg
T5-small Baseline -5.93 -5.39 -3.73 -5.02
Ours -6.81 -6.95 -6.39 -6.72
FT5-small Baseline -2.12 -5.07 -1.69 -2.96
Ours -2.49 -4.95 -2.66 -3.37
T5-base Baseline 0.72 -2.45 -3.56 -1.76
Ours 0.93 1.57 -0.02 0.83
FT5-base Baseline -2.46 -2.45 -2.23 -2.38
Ours -1.94 -2.07 -3.58 -2.53
T5-large Baseline -3.41 -0.69 -0.02 -1.37
Ours -4.05 -4.53 -4.03 -4.20
FT5-large Baseline -1.01 0.20 -0.16 -0.32
Ours -0.26 0.71 -0.23 0.07
(A) Baseline
(B) Ours
(c) B-A
Figure 4: Heatmaps of backward transfer scores on previous tasks for Order 5. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 5: Per-task BWT comparison between our method (blue) and the baseline (red) for Order 5. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 6: Heatmaps of backward transfer scores on previous tasks for Order 6. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 7: Per-task BWT comparison between our method (blue) and the baseline (red) for Order 6. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 8: Heatmaps of backward transfer scores on previous tasks for Order A. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 9: Per-task BWT comparison between our method (blue) and the baseline (red) for Order A. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 10: Heatmaps of backward transfer scores on previous tasks for Order B. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 11: Per-task BWT comparison between our method (blue) and the baseline (red) for Order B. Positive bars indicate improved retention of prior tasks.
(A) Baseline
(B) Ours
(c) B-A
Figure 12: Heatmaps of backward transfer scores on previous tasks for Order C. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
Figure 13: Per-task BWT comparison between our method (blue) and the baseline (red) for Order C. Positive bars indicate improved retention of prior tasks.
Table 20: Summary of the results with the T5-large model on the short-sequence benchmark. Average accuracy after training on the last task is reported. All results are averaged over 3 runs. For the T5 model, we follow few-shot CL settings as in [25]
Method Order 1 Order 2 Order 3 avg
Per-task $70.0$ $70.0$ $70.0$ $70.0$
Continual Finetune $18.9$ $24.9$ $41.7$ $28.5$
Data Replay $35.4$ $37.1$ $41.5$ $38.0$
EWC $39.0$ $38.0$ $44.8$ $40.6$
LFPT5 $47.6$ $52.6$ $57.9$ $52.7$
CODA-Prompt $77.0$ $80.2$ $75.9$ $77.7$
ProgPrompt $74.1$ $74.2$ $75.3$ $74.5$
SHLPT $80.2$ $79.7$ $80.9$ $80.1$
Ours* 74.84 74.41 74.55 74.60
B.3 Implementation Details
In our experiments, we focus on encoder-decoder transformer-based architecture, adopting both the original T5[26] and Flan-T5[7] as our backbone architectures. To study the impact of model size, we perform evaluations using the small, base, and large variants. This setup allows us to assess the scalability of our method across a range of model capacities.
Experimental Setting
In all our experiments with T5 and Flan-T5 backbones, we fix the prompt length to 10 tokens, train for 10 epochs, and use 1k representative samples per class. [^3] We could not experiment with 300 epochs and 20 samples per class (as in ProgPrompt [27]) due to resource constraints. This configuration is applied uniformly across the small, base, and large variants of both model families, ensuring a fair yet efficient comparison to prior work such as ProgPrompt [27]. All experiments were performed on a single A100 40GB GPU with batch size set to 8 for all experiments.
Table 21: Examples highlighting inconsistent label mappings in COPA under task-agnostic inference. Prior works like Progressive Prompts use manual conversions (e.g., 0 $\rightarrow$ “false”), often misaligning task semantics. GRID mitigates these inconsistencies by inferring appropriate labels automatically.
Premise Choice1 Choice2 Question Label ProgPrompt GRID
The man lost the competition. The competition was sabotaged. He intimidated his competitors. cause 0 false choice1
I regained composure from my fit of anger. My heart pounded. I took deep breaths. cause 1 true choice2
The cook’s eyes watered. He ran out of onions. He cut an onion. cause 1 true choice2
The tree branch landed in the river. The branch moved downstream. The river’s current became stronger. effect 0 false choice1
The woman retired. She received her pension. She paid off her mortgage. effect 0 false choice1